OpenStudy (anonymous):

Need help solving this homogeneous equation. (xy+y^2)dx-x^2dy=0. This is what I did so far, dy/dx= (-xy-y^2)/(-x^2) Let z=(y/x) so multiply the right by (y/x^3) I got, ((-y/x)^2-(y/x)^3)/(-y/x). Is this right so far? As I went on things got confusing.

OpenStudy (amistre64):

if z = y/x, doesnt y = zx?

OpenStudy (amistre64):

(xy+y^2)dx-x^2dy=0 (x(zx)+(zx)^2)dx -x^2dy=0 (x(zx)+(zx)^2) -x^2 dy/dx=0 well, since zx = y, then z'x + z = y', or =dy/dx (x(zx)+(zx)^2) -x^2(z'x + z)=0 but then its still not a linear equation is it, since we have a z^2 instead of a y^2

OpenStudy (amistre64):

the sub of z = y/x is useful if it makes the setup seperable, does this make it seperable?

OpenStudy (anonymous):

it does. After I get dy/dx into the form of (y/x) I can then substitute z and rewrite my equation as x(dz/dx)+z. So I get, x(dz/dx) = (z^2-z^3)/z - z

OpenStudy (anonymous):

does that make sense? I just learned this today so I may be wrong

OpenStudy (amistre64):

lets try this (xy+y^2)dx -x^2dy = 0 dividing by dx (xy+y^2) -x^2 dy/dx = 0 and rewriting in the y' notation (xy+y^2) -x^2 y' = 0 divide by x^2 y/x + (y/x)^2 -y' = 0 so z + z^2 - y' = 0 with the sub but y' = z'x + z z + z^2 -z'x - z = 0 z^2 -z'x = 0 ok, its looking better to me now

OpenStudy (dumbcow):

^^ yeah thats what i was getting.... now its separable:)

OpenStudy (anonymous):

that is making much more sense!! I am having trouble getting the equation into the form of (y/x). Like figuring out what to multiply by to get in that form

OpenStudy (amistre64):

its still not linear, but seperable doesnt care much about linearity :)

OpenStudy (amistre64):

dividing off the x^2 seemed reasonable to me so i tried it to see what it would do, it was a good call apparently :)

OpenStudy (anonymous):

okay! Thank you! I will try this out and see if it works :)

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

It worked!